LonePieceOfToast:

Jeremy accidently dropped his toy stuffed animal from the balcony of his apartment on the fourth floor. The toy hit the ground at a velocity of 16.0 meters/second. At impact, it took 2.0 seconds for the toy’s velocity to reach 0 meters/second. If the toy has a mass of 0.25 kilograms, what's the force of the impact? Use F = ma, where . The force of the impact is _____ newtons.

lowkey:

F = (mass)(acceleration) = ma m = 0.25 kg Vi = 16 m/s t = 2 s Vf = 0 m/s (since it was put to stop) a=(Vf-Vi)/t a=(0-16)/2 a = 8 m/s^2 (decelerating) F = ma = (0.25 kg)(8 m/s^2) F = 2 N